Hello, can someone please help?
Condense the following:
1. log3 2x+3log3 4x
2. ln72x-2 ln 2 y
Use the Chane of base formula
Log4 31
log7 100

Answer by stanbon(48535)

Condense the following:
1. log3 2x+3log3 4x
= log3 [ 2x*(4x)^3]
= log3 [ 2x*64x^3]
= log3 [128x^4]
=================

2. ln72x-2 ln 2 y
= ln(72x) - ln(2y)2
= ln(72x) - ln(4y^2)
= ln[72x/4y^2]
= ln[18x/y^2]
===================
Use the Change of base formula

Log4 31
= log(31)/log(4)
= 2.477
=======================
log7 100
= log(100)/log(7)
= 2/0.8451
= 2.3666
==================
Cheers,
Stan H.